# Hidden Markov Model Filtering with Equal Exit Probabilities

### v1

@@ II Steady State Dynamical Analysis
where $\delta$ is the step-size parameter. Equations (18) and (19) both apply a discount to the information from the previous time step during the iteration process. The subtle difference between them lies in the fact that the latter normalizes the weighting between past and new information through the step-size parameter $\delta$.
It can be seen that the recursion (10) is both nonlinear and stochasticnonlinear since the first term of the r.h.s. is a nonlinear function of $\boldsymbol{x}_{m,i-1}$ and stochastic since the system is driven by the function of a random process $\boldsymbol{\xi}_{i}$.
To facilitate analysis, we introduce the following deterministic reference dynamical system, where stochastic quantities are replaced by their expected values:
$\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-d_{m},$ (20)
where $m=1,\ldots.M-1$ and
$d_{m}$ $\triangleq D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{m}))-D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{0}))$ (21)
$=-\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}\right].$
Assumption 2 guarantees that $d_{m}>0$. From the definition of $d_{m}$, it is evident that it characterizes the agent’s ability to distinguish any incorrect state $\theta_{m}$ from the underlying true state $\theta_{0}$ using the likelihood model and the observation $\boldsymbol{\xi}_{i}$. A larger $d_{m}$ indicates stronger identifiability.
In this paper, we first analyze the $(M-1)$-dimensional deterministic dynamical system given by (20) as a reference system. Subsequently, we will demonstrate that due to the contraction properties of the function $F(\cdot)$, the asymptotic difference between the original stochastic system (10) and this reference system can be bounded.
In the following theoretical analysis, we focus on the scenario where $0<\alpha<1/M$, ensuring $10<1-\alpha M<$ 1. This condition, referred to as slow-changing conditions, not only aligns better with intuitive expectations—since a high value of $\alpha$ (indicative of highly volatile environments) is expected to degrade the algorithm’s performance—but also facilitates a more analytically tractable analysis of the algorithm.
II-A Convergence Rate and Adaption Ability
In this part, we will prove that the reference dynamical system (20) will converge to a fixed point at an exponential rate. Based on this, we can analyze the adaptation capability and how it can be influenced by the exit probability parameter $\alpha$ and $d_{m}$.
The following theorem provides the main result regarding the convergence of the reference dynamical system (20) and its convergence rate.
Theorem 1
When $0<\alpha<1/M$, $\hat{x}_{i}$ defined in (20) converges to a unique fixed point, i.e.,
$\lim_{i\rightarrow\infty}\hat{x}_{i}=\hat{x}^{\infty},$ (22)
@@ II-A Convergence Rate and Adaption Ability
$\hat{x}_{i}=\left[\hat{x}_{1,i},\ldots,\hat{x}_{M-1,i}\right]^{\top},$ (23)
the fixed point is given by
$\hat{x}^{\infty}_{m}=\log\frac{\alpha}{\alpha\exp(d_{m})+(1-\alpha M)(\exp(d_{m})-1)\hat{\mu}_{0}^{\infty}},$ (25)
@@
Furthermore, the convergence rate is bounded by:
@@
where the contraction factor is given by:
$\lambda=1-\min\left\{\alpha,\frac{\underline{d}}{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\underline{d}}\right\},$ (28)
@@
Proof:
Proof omitted due to space limitations. ∎
Remark 2
Theorem 1 shows that the reference dynamical system converges to a fixed point, which is strictly less than zero for all components. To see this, for all $m=1,\ldots,M-1$, when $01-\alpha M>$ 0 we have:
$\alpha\exp(d_{m})+(1-\alpha M)(\exp(d_{m})-1)\hat{\mu}^{\infty}_{0}>\alpha\exp(d_{m}),$ (31)
implying
$x_{m}^{\infty}<-d_{m}<0.$ (32)
This result means that in the reference dynamical system, the agent places higher belief on the underlying true state at the fixed point. When $01-\alpha M>$ 0, the higher the value of $d_{m}$, which represents the agent’s ability to distinguish between $\theta_{0}$ and $\theta_{m}$, the more confident the agent is in the underlying true state at the fixed point.
Remark 3
Theorem 1 provides an upper bound on the convergence rate of $\hat{x}_{i}$ to $\hat{x}^{\infty}$ in the sense of the $L_{\infty}$-norm, which decreases as the exit probability $\alpha$ increases, as shown in (28). We can also observe the impact of the data quality, captured in $\underline{d}$, and the initial setting, captured in $\overline{x}_{0}$, on the convergence rate of the reference dynamical system to its fixed point. The parameter $\underline{d}$ depends on both the local likelihood function (e.g., pre-trained models) and the quality of observations (e.g., variance of local data noise). This parameter, referred to as the identifiability parameter, reflects the agent’s ability to distinguish between the two most challenging states. From Theorem 1, it follows that higher data quality—manifested as a larger $\underline{d}$—accelerates the convergence to the fixed point, while the deviation in the initial belief, captured in $\overline{x}_{0}$, slows down the convergence rate.
The impact of the exit probability $\alpha$ and the identifiability parameter $\underline{d}$ on the convergence rate can be further illustrated by the following lemma.
Corollary 1
When $0<\alpha<1/M$, the convergence rate of the reference dynamical system has the following property:
@@
$\gamma=\frac{(1-\alpha M)\left(1+2\alpha U(M-1)\right)}{(1-\alpha M+\alpha)^{2}}\exp(-\underline{d}),$ (34)
@@
then a preciser bound of contraction rate can be given by:
$\left\|\exp(\hat{x}_{i})-\exp(\hat{x}^{\infty})\right\|_{1}\leq\gamma^{\prime}\left\|\exp(\hat{x}_{i-1})-\exp(\hat{x}^{\infty})\right\|_{1},$ (38)
@@
$\gamma^{\prime}=\frac{\exp(-\underline{d})(1-\alpha M)}{(1-\alpha M+\alpha)^{2}}.$ (39)
Proof:
Proof omitted due to space limitations. ∎
Corollary 1 further shows that as $\alpha\rightarrow 1/M$, contraction coefficient $\gamma\rightarrow 0$, indicating that the rate of convergence accelerates. This observation aligns with intuition: when $\alpha\rightarrow 1/M$ in equation (6), the system relies less on historical observations and becomes more reliant on the most recently observed data. Additionally, as $\underline{d}\rightarrow\infty$, $\gamma\rightarrow 0$, which again shows that higher data quality leads to a faster convergence rate of the system to its fixed point.
@@
The convergence rate of the reference dynamical system to its fixed point, as presented in Theorem 1, plays a crucial role in understanding the adaptation capabilities of the algorithm. Specifically, when the environment remains in a fixed state for an extended period, the reference dynamical system converges to its fixed point at an exponential rate. When the environment state changes, the corresponding fixed point shifts accordingly, and the system will converge to the new fixed point exponentially fast, as illustrated in Fig. 1. The convergence speed depends on the choice of the parameter $\alpha$ and the quality of the data and model. When using the traditional Bayes’ formula (3), as indicated by equations (11) and (20) with $\alpha=0$, the corresponding deterministic dynamical system for the log-belief ratio reduces to a linear equation, which evolves linearly over time. This slower rate explains why the proposed algorithm exhibits superior adaptability in dynamic environments.
@@
Figure 1: (a) The agent decides between two possible states, $\theta_{0}$(state 1) and $\theta_{1}$(state 2), where the likelihood models for these two states, as well as the model for environmental observations under these two states, are both standard normal density functions with distinct means. The environment remains in state 1 until time $T_{1}$, after which it switches to state 2 for the remaining duration. (b) The log-belief ratio $\log(\boldsymbol{\mu}_{i}(\theta_{1})/\boldsymbol{\mu}_{i}(\theta_{0}))$ is compared over time for both the $\alpha$-HMM and the traditional Bayes’ filter. Here $T^{A}_{\text{adapt}}$ and $T^{B}_{\text{adapt}}$ represent the adaptation time—specifically, the time required for the log-belief ratio to shift from negative to positive—required for the $\alpha$-HMM and Bayes’ formula, respectively. The HMM filter quickly converges and stabilizes at the fixed point, whereas the Bayes’ filter continues to accumulate ”confidence” in the current state. When the environment changes, the Bayes’ filter must counteract the previously accumulated confidence and adjust at a slower, linear rate. This slower adjustment prevents the Bayes’ filter from rapidly responding to changes in the true state. The difference in adaptability between the two algorithms is particularly pronounced when $T_{1}$ is large. (c) The belief on the true state under both algorithms over time. In deterministic expected dynamics, the $\alpha$-HMM exhibits superior adaptability, with the belief remaining above 0.5 for most of the time, indicating correct decision-making. When observation noise is present, the $\alpha-$HMM algorithm exhibits larger fluctuations to maintain adaptability. This trade-off allows the $\alpha$-HMM to remain responsive, albeit at the cost of increased variance in belief estimates.
II-B Fixed Point Value and Learning Performance

### v2

@@ 2 Steady State Dynamical Analysis
where $\delta$ is the step-size parameter. Equations (33) and (32) both apply a discount to the information from the previous time step during the iteration process. The subtle difference between them is that the former employs two independent parameters, where $\alpha$ characterizes the rate of change of the environment, and $\beta$ implements the tradeoff between memory and new information. As remarked earlier, this added flexibility is useful because the choice $\beta=1$ recovers the optimal HMM-filter for a Markov chain with equal exit probability. Setting $\alpha=\frac{\delta}{M}$, $\beta=\delta$ instead yields adaptive social learning variant from [28]. We will illustrate in the numerical results how different parameter choices can yield improved performance in specific settings.
2.1 Convergence Rate and Adaption Ability
In this part, we will prove that the reference dynamical system (27) will converge to a fixed point at an exponential rate. Based on this, we can analyze the adaptation capability and how it can be influenced by the exit probability $\alpha$, step-size parameter $\beta$ and identifiability vector $d$ defined in (21).
Definition 1 (Contractions)
Let $(X,\rho)$ be a metric space. A mapping $T:X\to X$ is a contraction, if there exists a constant $c$, with $10\leq c<$ 1, such that
$\rho\left(T(x),T(y)\right)\leq c\rho(x,y)$ (34)
for all $x,y\in X$.
Based on the above definition, we now present the following lemma, which demonstrates that the mapping $F(x)$ defined in (23) is a contraction on the non-positive region; this property facilitates the convergence analysis of the reference dynamical system.
Lemma 1
Consider $0<\alpha<1/M$ and $\beta>0$, the mapping $F$ defined in (23) and (18) has the following properties:
  1. 1.
    $F(\left(-\infty,0\right]^{M-1})\subset\left(-\infty,0\right]^{M-1}$; that is, $F$ maps the non-positive region into itself.
  2. 2.
    $F$ is a contraction on $\left(-\infty,0\right]^{M-1}$ with respect to the $L_{\infty}-$norm.
Proof 2.1.
Appendix 7.
Let $T:X\to X$ be a mapping. A point of $x\in X$ such that
$T(x)=x$ (35)
is called a fixed point of $X$. For a discrete dynamical system $x_{i}=T(x_{i-1})$, a fixed point of the map $T$ corresponds to an equilibrium state of the system. In the remainder of this paper, we will consistently refer to it as the fixed point.
For the reference dynamical system (27), its fixed point has the following property:
Lemma 2.2.
The fixed point $\hat{x}^{\infty}$ of the reference dynamical system (27) satisfies:
$\hat{x}^{\infty}_{m}=\log\frac{\alpha}{\alpha\exp(\beta d_{m})+(1-\alpha M)(\exp(\beta d_{m})-1)\hat{\mu}_{0}^{\infty}},$ (37)
@@ 2.1 Convergence Rate and Adaption Ability
Proof 2.3.
Supplementary Material I.
Lemma 2.2 shows that the reference dynamical system has a fixed point, which is strictly negative for all components. To verify this, for all $m=1,\ldots,M-1$, when $01-\alpha M>$ 0 we have:
$\alpha\exp(\beta d_{m})+(1-\alpha M)(\exp(\beta d_{m})-1)\hat{\mu}^{\infty}_{0}>\alpha\exp(\beta d_{m}),$ (39)
implying
$\hat{x}_{m}^{\infty}<-\beta d_{m}<0.$ (40)
This result implies that the reference dynamical system learns correctly, since the log-belief ratio $\hat{x}_{m,i}$ is strictly less than zero for any wrong state $\theta_{m}$ When $01-\alpha M>$ 0, increasing either $\beta$ (the weight assigned to the observational data) or $d_{m}$ (the distinguishability between $\theta_{0}$ and $\theta_{m}$) drives the fixed-point value $\hat{x}_{m}^{\infty}$ further negative, indicating stronger confidence in the true state relative to the competing state $\theta_{m}$.
Now we are able to provide the main result regarding the convergence of the reference dynamical system (27) and its convergence rate.
Theorem 2.4.
When $0<\alpha<1/M$ and $\beta>0$, $\hat{x}_{i}$ defined in (27) will converge to a unique fixed point, i.e.,
$\lim_{i\to\infty}\hat{x}_{i}=\hat{x}^{\infty},$ (41)
and
@@
where
$\lambda=1-\min\left\{\frac{2\alpha}{1-\alpha M+2\alpha},\frac{\beta\underline{d}}{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\beta\underline{d}}\right\},$ (43)
@@
Proof 2.5.
Appendix 8
Remark 2.6.
Theorem 2.4 provides an upper bound on the convergence rate of $\hat{x}_{i}$ to $\hat{x}^{\infty}$ in the sense of the $L_{\infty}$-norm, which decreases as the exit probability $\alpha$ increases, as shown in (43). We can also observe the impact of the data, captured in $\beta\underline{d}$, and the initial setting, captured in $\overline{x}_{0}$, on the convergence rate of the reference dynamical system to its fixed point. The parameter $\underline{d}$ depends on both the local likelihood function (e.g., pre-trained models) and the quality of observations (e.g., variance of local data noise). This parameter, referred to as the identifiability parameter, reflects the ability to distinguish between the two most challenging states. From Theorem 2.4, it follows that higher identifiability $\underline{d}$ and higher $\beta$ accelerate the convergence to the fixed point, while the deviation in the initial setting, captured in $\overline{x}_{0}$, slows down the convergence rate.
The impact of the exit probability $\alpha$, step-size $\beta$ and the identifiability parameter $\underline{d}$ on the convergence rate are further illustrated by the following corollary.
Corollary 2.7.
When $0<\alpha<1/M$ and $\beta>0$, the convergence rate of the reference dynamical system is characterized by the following inequality, which further reveals the influence of the identifiability parameter $\underline{d}$:
@@
$\gamma=\frac{(1-\alpha M)\left(1+2\alpha U(M-1)\right)}{(1-\alpha M+\alpha)^{2}}\exp(-\beta\underline{d}),$ (47)
@@
then a more precise bound of the contraction rate can be given by:
$\left\|\exp(\hat{x}_{i})-\exp(\hat{x}^{\infty})\right\|_{1}\leq\gamma_{1}\left\|\exp(\hat{x}_{i-1})-\exp(\hat{x}^{\infty})\right\|_{1},$ (51)
@@
$\gamma_{1}=\frac{\exp(-\beta\underline{d})(1-\alpha M)}{(1-\alpha M+\alpha)^{2}}.$ (52)
Proof 2.8.
Supplementary Material II.
Corollary 2.7 provides a more precise upper bound on the convergence rate of the reference dynamical system to its fixed point. It shows that as $\alpha\to 1/M$, the contraction coefficient $\gamma\to 0$, indicating that the rate of convergence accelerates. This observation aligns with intuition: when $\alpha\to 1/M$ in equation (5), the system relies less on historical observations and becomes more reliant on the most recently observed data. In subsequent numerical experiments, we observe that, generally larger values of $\alpha<1/M$ lead to faster convergence. Additionally, as $\beta\to\infty$, $\gamma\to 0$, which shows that increasing the value of $\beta$ leads to a faster convergence rate of the system to its fixed point.
@@
The convergence rate of the reference dynamical system to its fixed point, as presented in Theorem 2.4, plays a crucial role in understanding the adaptation capabilities of the algorithm. Specifically, when the environment remains in a fixed state for an extended period, the reference dynamical system converges to its fixed point at an exponential rate. When the environment state changes, the corresponding fixed point shifts accordingly, and the system will converge to the new fixed point exponentially fast, as illustrated in Fig. 2. The convergence speed depends on the choice of the parameters $\alpha$ and $\beta$, and the quality of the data and model. When using the traditional Bayesian update (6), as indicated by equations (17) and (27) with $\alpha=0$ and $\beta=1$, the corresponding deterministic dynamical system simplifies to the linear equation
$\hat{x}_{i}=\hat{x}_{i-1}-d,$ (54)
which shows that the log-belief ratio evolves linearly over time. This slower rate explains why the proposed algorithm exhibits superior adaptability in dynamic environments.
Example 2.10.
To further illustrate the adaptability difference between $\alpha\beta$-HMM and the Bayesian update (6), consider a binary decision-making problem with candidate states, $\theta_{0}$ and $\theta_{1}$. Before time instant $T_{1}$, the environmental signals are more aligned with the likelihood function corresponding to $\theta_{0}$, whereas after $T_{1}$, they become more aligned with the likelihood function corresponding to $\theta_{1}$.
We analyze the expected evolution of the log-belief ratio:
$\boldsymbol{x}_{i}=\log\frac{\boldsymbol{\mu}_{i}(\theta_{1})}{\boldsymbol{\mu}_{i}(\theta_{0})}.$ (55)
For the traditional Bayes’ formula (6), we have:
$\hat{x}_{i}=\hat{x}_{i-1}+\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{1})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}\right]=\begin{cases}\hat{x}_{i-1}-d_{\text{s}1}&\text{if }i\leq T_{1},\\ \hat{x}_{i-1}+d_{\text{s}2}&\text{otherwise},\end{cases}$ (56)
where
$d_{\text{s}1}=D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{1}))-D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{0}))>0,$ (57)
and
$d_{\text{s}2}=D_{\text{KL}}(f(\cdot|\theta_{1})||L(\cdot|\theta_{0}))-D_{\text{KL}}(f(\cdot|\theta_{1})||L(\cdot|\theta_{1}))>0.$ (58)
Starting from an initial value $\hat{x}_{0}$, $\hat{x}_{i}$ will initially decrease linearly with time before $T_{1}$ and then increases. Let $T_{\text{adapt}}$ denote the time required for correcting past beliefs, that is, the time needed for adaptation to the new environment. Specifically, $\hat{x}_{i}>0$ when $i>T_{1}+T_{\text{adapt}}$. For the Bayesian update(6), we have:
$T_{\text{adapt}}>\frac{d_{\text{s}1}T_{1}-\hat{x}_{0}}{d_{\text{s}2}}.$ (59)
For $\alpha\beta$-HMM, we have:
$\hat{x}_{i}=F(\hat{x}_{i-1})+\beta\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{1})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}\right],$ (60)
Following Theorem 2.4, we can see that:
$\left|\hat{x}_{T_{1}}-\hat{x}_{\text{s}1}^{\infty}\right|\leq\lambda_{\text{s}1}^{T_{1}}\left|\hat{x}_{0}-\hat{x}_{\text{s}1}^{\infty}\right|,$ (61)
and when $i>T_{1}$,
$\left|\hat{x}_{i}-\hat{x}_{\text{s}2}^{\infty}\right|\leq\lambda_{\text{s}2}^{i-T_{1}}\left|\hat{x}_{T_{1}}-\hat{x}_{\text{s}2}^{\infty}\right|,$ (62)
where $\hat{x}_{\text{s}1}^{\infty}<0$ and $\hat{x}_{\text{s}2}^{\infty}>0$ are the two fixed points before and after $T_{1}$, respectively, and $\lambda_{\text{s}1}$ and $\lambda_{\text{s}2}$ are the convergence coefficients in two stages defined as in Theorem 2.4.
The adaptation time $T_{\text{adapt}}$ for $\alpha\beta$-HMM should satisfy:
$\hat{x}_{T_{1}+T_{\text{adapt}}}>0.$ (63)
A sufficient condition for (63) is:
$\left|\hat{x}_{T_{1}+T_{\text{adapt}}}-\hat{x}_{\text{s}2}^{\infty}\right|<\left|\hat{x}_{\text{s}2}^{\infty}\right|.$ (64)
Following (61) and (62), we have for $i>T_{1}$,
$\left|\hat{x}_{i}-\hat{x}_{\text{s}2}^{\infty}\right|$ $\leq\lambda_{\text{s}2}^{i-T_{1}}\left|\hat{x}_{T_{1}}-\hat{x}_{\text{s}2}^{\infty}\right|$ (65)
$\leq\lambda_{\text{s}2}^{i-T_{1}}\left(\left|\hat{x}_{T_{1}}-\hat{x}_{\text{s}1}^{\infty}\right|+\left|\hat{x}_{\text{s}2}^{\infty}-\hat{x}_{\text{s}1}^{\infty}\right|\right)$
$\leq\lambda_{\text{s}2}^{i-T_{1}}\left(\lambda_{\text{s}1}^{T_{1}}\left|\hat{x}_{0}-\hat{x}_{\text{s}1}^{\infty}\right|+\left|\hat{x}_{\text{s}2}^{\infty}-\hat{x}_{\text{s}1}^{\infty}\right|\right).$
To guarantee (64), it is sufficient that:
$\lambda_{\text{s}2}^{T_{\text{adapt}}}\left(\lambda_{\text{s}1}^{T_{1}}\left|\hat{x}_{0}-\hat{x}_{\text{s}1}^{\infty}\right|+\left|\hat{x}_{\text{s}2}^{\infty}-\hat{x}_{\text{s}1}^{\infty}\right|\right)<\left|\hat{x}_{2}^{\infty}\right|,$ (66)
which implies
$T_{\text{adapt}}>\frac{\log\frac{\left|\hat{x}_{2}^{\infty}\right|}{\lambda_{\text{s}1}^{T_{1}}\left|\hat{x}_{0}-\hat{x}_{\text{s}1}^{\infty}\right|+\left|\hat{x}_{\text{s}2}^{\infty}-\hat{x}_{\text{s}1}^{\infty}\right|}}{\log\lambda_{\text{s}2}}.$ (67)
Comparing (59) and (67), it can be observed that in Bayes’ formula (6), the adaptation time increases linearly with $T_{1}$, and is highly dependent on the initial value $\hat{x}_{0}$. In contrast, under the $\alpha\beta$-HMM strategy, the influence of the initial value $\hat{x}_{0}$ vanishes as $T_{1}$ approaches infinity. Furthermore, since $\lambda_{\text{s}1}<1$, the adaptation time will not increase indefinitely with $T_{1}$, but instead will stabilize at a certain value. It is worth noting that, under similar settings, [28] also derives an adaptation time in the case of a single agent, which also exhibits the independence of the previous stationary interval $T_{1}$ when it is sufficiently large. Figure 2 presents a comparison of the adaptability of Bayes’ formula and $\alpha\beta$-HMM under different $T_{1}$ values in this example.
@@
Figure 2: (a) Two candidate states are considered, $\theta_{0}$(state 1) and $\theta_{1}$(state 2), where the likelihood models for these two states, as well as the model for environmental observations under these two states, are both standard normal density functions with distinct means. The environment remains in state 1 until time $T_{1}$, after which it switches to state 2 for the remaining duration. (b) The log-belief ratio $\log(\boldsymbol{\mu}_{i}(\theta_{1})/\boldsymbol{\mu}_{i}(\theta_{0}))$ is compared over time for both the $\alpha\beta$-HMM and the traditional Bayes’ filter (6). Here $T^{A}_{\text{adapt}}$ and $T^{B}_{\text{adapt}}$ represent the adaptation time required for the $\alpha\beta$-HMM and Bayes’ formula, respectively. The $\alpha\beta$-HMM filter quickly converges and stabilizes at the fixed point, whereas the Bayes’ filter continues to accumulate ”confidence” in the current state. When the environment changes, the Bayes’ filter must counteract the previously accumulated confidence and adjust at a slower, linear rate. This slower adjustment prevents the Bayes’ filter from rapidly responding to changes in the true state. The difference in adaptability between the two algorithms is particularly pronounced when $T_{1}$ is large. (c) The belief on the true state under both algorithms over time. In deterministic expected dynamics, the $\alpha\beta$-HMM exhibits superior adaptability, with the belief remaining above 0.5 for most of the time, indicating correct decision-making. When observation noise is present, the $\alpha\beta$-HMM algorithm exhibits larger fluctuations to maintain adaptability. This trade-off allows the $\alpha\beta$-HMM to remain responsive, albeit at the cost of increased variance in belief estimates.
2.2 Fixed Point Value and Learning Performance